Suppose you obtain a 25-year mortgage loan of $199,000 at an annual interest rate of 8.2%. The annual property tax bill is $974 and the annual fire insurance premium is $487. Find the total monthly payment for the mortgage, property tax, and fire insurance. (Round your answer to the nearest cent.)
Answer #1

А 1 Monthly loan payment 2 Property tax bill 3 Fire insurance premium 1,562.37 =PMT(8.2%/12,25*12,-199000,0,0) 81.17 =974/12

*Please rate thumbs up